# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v4

@@ Abstract
The optimal operation of transportation systems is often susceptible to unexpected disruptions, such as traffic accidents and social events. Many established control strategies relying on mathematical models can struggle with real-world disruptions, leading to a significant deviation from their anticipated efficiency. This work applies the cutting-edge concept of antifragility to design a traffic control strategy for urban road networks against disruptions. Antifragility sets itself apart from robustness, resilience, and reliability as it represents a system’s ability to not only withstand stressors, shocks, and volatility but also to thrive and enhance performance in the presence of such adversarial events. Incorporating antifragile terms composed of traffic state derivatives and redundancy, a model-free deep reinforcement learning algorithm is developed and subsequently evaluated in a two-region cordon-shaped urban traffic perimeter network. Promising results highlight (a) the superior performance of the proposed algorithm compared to the state-of-the-art methods under incremental magnitude of disruptions, (b) distribution skewness as the antifragility indicator demonstrating its relative antifragility, and (c) its effectiveness under limited observability due to real-world data availability constraints. The proposed antifragile methodology is generalizable and holds potential for application beyond perimeter control, offering integration into systems exposed to disruptions across various disciplines.
@@
journal: Elsevier
@@ 1 Introduction
Transportation networks serve as vital channels for the movement of people and the flow of goods, and the optimization and control of intelligent transportation systems have led to a multitude of research endeavors and practical implementations (Auer et al., 2016). Given that various sorts of disruptions, such as traffic accidents, social events, and adversarial weather conditions, often occur unexpectedly in real-world networks, examining the robustness and resilience of transportation systems is highly crucial (Ganin et al., 2019). What makes it even more challenging is the continuous growth of motorized traffic. For instance, an approximate 50% increase in traffic demand can be observed in both the U.S. (U.S. Department of Transportation, 2019) and Switzerland (Federal Statistical Office, 2020) over the past few decades. Such growth can lead to not only an escalation of congestion but also more traffic accidents (Dickerson et al., 2000). To picture the consequences, recent work in Sun et al. (2024) has demonstrated the fragile nature of road transportation networks with mathematical proof, demonstrating that performance degrades exponentially with a linearly increasing magnitude of disruptions. Therefore, transportation systems must be designed to secure a decent level of service even when faced with unexpected disruptions of unforeseen magnitudes and the exponentially growing negative effects.
@@
To address such issues, the concept of antifragility has shed light on a potential solution. First introduced in the reputable general publication Antifragile: Things That Gain from Disorder (Taleb, 2012), and later mathematically formulated as academic works in Taleb (2013); Taleb and Douady (2013), antifragility has provided insights into designing systems that can benefit from disruptions and perform better under growing volatility and randomness. Its counterpart concept, fragility, can be dated earlier in complexity theory in Vespignani (2010), indicating a cascading effect of interdependent variables in complex networks, such as in transportation networks (Cats and Hijner, 2021). Ever since the concept was proposed, antifragility has gained substantial interest from both the public and academia, particularly in the complex systems (Axenie et al., 2024) and risk engineering communities (Aven, 2015; Thekdi and Aven, 2019; Grassi et al., 2024). However, the design principles for realizing antifragile transportation systems are generally unexplored. One promising approach to induce antifragility is applying learning-based algorithms, such as Reinforcement Learning (RL) (Haydari and Yılmaz, 2022), since RL agents can gradually adjust their decision-making when deployed to an environment subject to variations.
The main goal of this paper is to design an antifragile perimeter control algorithm capable of adapting to increasing magnitudes of disruption, reflecting the effects of urban densification and rising traffic demand. The key contributions are: a) We formulate and distinguish the concept of antifragility as opposed to other related and commonly used terms in the transportation domain; b) We introduce how antifragility can be incorporated into RL algorithms to achieve superior performance compared to benchmark methods, tackling both fragile performance and observability issues; and c) We adopt a skewness-based quantitative indicator to showcase the antifragile properties of our proposed algorithm under increasing disruptions. A contribution map following previous works is illustrated in Fig. 1.
@@ 2 Literature review
This section reviews the relevant literature on three topics intertwined in this work. First, a macroscopic traffic model and the control strategy applied in this paper are introduced, which serve as the basis of model dynamics for the simulation environment. Next, we study how robustness or resilience can be induced with RL algorithms to counter real-world model uncertainty. Finally, related terminologies to antifragility are distinguished while introducing the design philosophy of an antifragile system.
@@ 2.1 Macroscopic fundamental diagram and perimeter control
Alleviating urban network congestion can be realized through various traffic control strategies. Since Geroliminis and Daganzo (2008) demonstrated the presence of the Macroscopic Fundamental Diagram (MFD) with empirical data and Daganzo and Geroliminis (2008) generated the first MFD analytically with variational theory for a homogeneous urban area, the mathematical relationship between flow-density-speed has paved the path for the development of control strategies at a macroscopic level, enabling more computationally feasible real-time traffic control strategies for large-scale networks (Knoop et al., 2012), such as perimeter control (Keyvan-Ekbatani et al., 2012; Geroliminis et al., 2013; Kouvelas et al., 2017; Yang et al., 2017), congestion pricing (Zheng et al., 2012; Zheng and Geroliminis, 2016; Genser and Kouvelas, 2022), route guidance (Yildirimoglu et al., 2015; Fu et al., 2022).
@@
Perimeter control is among the strategies that have attracted immense attention and research efforts. By mitigating incoming flows from adjacent regions into a protected zone, the traffic density in the protected area remains below the critical density to uphold a satisfactory serviceability. Geroliminis et al. (2013) proposed an optimal perimeter control method using Model Predictive Control (MPC) and proved its effectiveness compared to a greedy controller in a cordon network. Later, Sirmatel and Geroliminis (2020) introduced a Moving Horizon Estimation (MHE) scheme together with MPC to tackle further measurement noise and the observability issues. One major issue of the previous works is the assumption of MFD homogeneity, and a substantial amount of effort has been made in investigating the partitioning algorithms so that a well-defined MFD can be created for each sub-network (Ambühl et al., 2019; Saedi et al., 2020), creating a multi-reservoir system for applying perimeter control (Aboudolas and Geroliminis, 2013). However, MFDs in the real world can hardly be well-defined, as demonstrated in Ambühl et al. (2021) with loop detector data over a year. Wang et al. (2015) and Ji et al. (2015) also showed that adverse weather conditions and traffic incidents can alter the shape of the MFDs, and even a recovery from the peak-hour congestion may lead to a hysteresis (Gayah and Daganzo, 2011). These phenomena could potentially violate the mathematical model that serves as the foundation for the established model-based perimeter controllers. Therefore, Zhou and Gayah (2021) introduced an RL-based algorithm that exhibits superior performance and offers the advantage of being model-free.
2.2 Leveraging RL to induce robustness or resilience
@@
Given the presence of real-world disturbances and disruptions, it is essential to assess the robustness and resilience of any newly developed traffic control algorithm. As RL offers flexibility in defining state representations, actions, and rewards, researchers have leveraged these features to enhance the robustness and resilience of RL-based traffic control strategies. Here, we present a collection of studies that apply RL in traffic control and claim to demonstrate robustness or resilience. While Zhou and Gayah (2021); Chen et al. (2022); Su et al. (2023); Zhou and Gayah (2023) focus on perimeter control frameworks, Aslani et al. (2018); Rodrigues and Azevedo (2019); Chu et al. (2020); Tan et al. (2020); Wu et al. (2020); Korecki et al. (2023) explore the design for signalized intersections. Some of these works directly compare the performance of the proposed algorithm with the benchmark methods under disruptions to showcase either robustness or resilience, whereas Rodrigues and Azevedo (2019); Tan et al. (2020); Chu et al. (2020); Zhou and Gayah (2023) further mention explicitly how it is induced with modifications of the algorithm. For instance, Rodrigues and Azevedo (2019) induced robustness by adding the elapsed time since the last green signal for each phase, Tan et al. (2020) experimented with speed or residual queue as an additional state representation in the RL algorithm, Chu et al. (2020) supplemented the control policies of neighboring intersections as additional information to the agents, and Zhou and Gayah (2023) used an extra binary congestion indicator in the state space. As a result, the algorithms are given additional information related to disturbances or disruptions of the environment. The analysis of the state-of-the-art RL studies considers reversals or sudden changes in the state-action-reward dynamics, which evoke unanticipated uncertainty. The problem in these contexts is often to respond to unexpected results appropriately, since they might indicate a shift in the environment. In this case, exploration refers to the process of looking for new information to improve the RL agent’s understanding of the traffic dynamics under disruptions, which would then be used to identify better courses of action. By exploring how robustness and resilience can be achieved using RL-based traffic control algorithms, we can potentially develop antifragile traffic control systems through similar approaches.
@@ 2.3 Antifragility: definitions and distinctions
Before diving into antifragile system design, we first need to review the distinctions between antifragility and other closely related terms, which include robustness, resilience, reliability, and adaptiveness. We follow the definitions proposed in Zhou et al. (2019), that robustness is concerned with assessing a system’s capacity to preserve its initial state and resist performance deterioration under minor disturbances, while resilience emphasizes the ability and speed of a system to recover from major disruptions to the original state. Reliability in transportation has a much broader range of meanings. Pennetti et al. (2020) proposed travel time reliability, which has a similar flavor to robustness in measuring deviation from normal operation, but focuses on the probability of such deviation exceeding a certain threshold. Cats et al. (2017) introduced a reliability assessment indicator to account for performance loss for the entire range of possible capacity reductions, whereas robustness pertains merely to a single disruption event. Adaptiveness describes the ability of systems to alter their traits to satisfy autonomy in different environments (Hooker, 2011), which aligns with the concept of proto-antifragility proposed in Taleb (2012). However, adaptiveness does not necessarily guarantee a consistent performance improvement under growing magnitudes of disruptions. Therefore, Taleb (2012) introduced the concept of antifragility, emphasizing the concave response of the system under increasing disruptions, which can be mathematically formulated with Jensen’s inequality $\mathbb{E}[g(X)]\leq g(\mathbb{E}[X])$. A graphical comparison between robustness, resilience, adaptiveness, and antifragility is shown in Fig. 16. It should be noted that while (anti-)fragility can be an innate property of a system, proper intervention and control strategies can lead the system to be more antifragile (Axenie et al., 2024). For instance, building on the mathematical proof by Sun et al. (2024) that urban road networks are naturally fragile, this paper seeks to develop antifragile solutions that mitigate such fragility through perimeter control.
@@
Ever since the concept of antifragility was proposed, it has become an increasingly popular concept across various disciplines in recent years, such as energy (Coppitters and Contino, 2023), electricity (Rachunok and Nateghi, 2020), biology (Kim et al., 2020), medicine (Axenie et al., 2022), cyber-systems (Chatterjee and Thekdi, 2020), and robotics (Axenie and Saveriano, 2023). Researchers have also proposed methods to incentivize the antifragile property of a system by emphasizing the derivatives of system states to capture the temporal evolution patterns of the system dynamics, i.e., how fast the system state deviates toward possible black swan events and the curvature of this deviation (Taleb and Douady, 2013; Taleb and West, 2023; Axenie et al., 2022). With this additional information, the system can anticipate ongoing disruptions and be more proactive under drastic changes. Similar to its function in resilience (Tan et al., 2019; Kamalahmadi et al., 2022), redundancy can also be added to the system to induce antifragility de Bruijn et al. (2020); Johnson and Gheorghe (2013); Munoz et al. (2022). Other feasible approaches also include time-scale separation and attractor dynamics (Axenie et al., 2022). However, leveraging the potential of antifragile system design for the operation of transportation networks is still a largely novel and unexplored notion.
@@ 3 Problem formulation
This paper studies perimeter control between two homogeneous cordon-shaped urban networks, as in Geroliminis et al. (2013), with the inner region representing a city center, as shown in Fig. 2(a). The total number of vehicles in region $i$ at time $t$ is denoted as $n_{i}(t)$, while the Origin-Destination (OD) pair can be further divided as $n_{ij}(t)$. The inner and outer regions are assumed to have different MFDs to represent the capacity difference of accommodating vehicles within the network between the city center and the surrounding region, which are defined as $G_{i}(n_{i}(t))$ as illustrated in Fig. 2(b). The total trip completion rate $M_{i}(t)$ for region $i$ at time $t$ can be determined through the corresponding MFD, which comprises both the intraregional trip completion $M_{ii}(t)$ and the interregional transfer flow $M_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. While $q_{i}$ denotes the new demand within region $i$ at time $t$, $q_{ij}(t)$ represents the additional demand based on an Origin-Destination (OD) pair from region $i$ to region $j$. To protect regions from being overflown by high traffic demand, the percentage of the transfer flow $M_{ij}(t)$ allowed to go across the perimeter at time $t$ is regulated by two perimeter controllers, denoted as $u_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. A list of all notations used in this paper, including the notations used in defining the RL algorithm and the antifragile terms, is summarized in Table 1.
@@
1. General notations in problem formulation
$t$ Time

### v5

@@ Abstract
The optimal operation of transportation systems is often susceptible to unexpected disruptions, such as traffic accidents and social events. Many established control strategies reliant on mathematical models can struggle with real-world disruptions, leading to significant divergence from their anticipated efficiency. This work fuses the cutting-edge concept of antifragility into the design of a learning-based traffic control strategy for large-scale urban road networks to perform against disruptions. Antifragility sets itself apart from robustness, resilience, and reliability as it represents a system’s ability to not only withstand stressors, shocks, and volatility but also to thrive and enhance performance in the presence of such adversarial events. Incorporating antifragile modules composed of traffic state derivatives and redundancy, a model-free deep reinforcement learning algorithm is developed and subsequently evaluated in a two-region cordon-shaped urban traffic perimeter network, and a numerical case study based on real-world data. Promising results highlight (1) the superior performance of the proposed algorithm compared to the state-of-the-art methods under incremental magnitude of disruptions, (2) distribution skewness as the antifragility indicator demonstrating its relative antifragility, (3) its effectiveness under limited observability due to real-world data availability constraints, and (4) robustness of the proposed antifragile module when embedded into advanced RL algorithms for superior performance. The proposed antifragile methodology is generalizable and holds potential for applications beyond perimeter control, offering integration into systems exposed to disruptions across various disciplines.
@@
journal: Reliability Engineering & System Safety
@@ 1 Introduction
Transportation networks serve as vital channels for the movement of people and the flow of goods, and the optimization and control of intelligent transportation systems have led to a multitude of research endeavors and practical implementations. Given that various sorts of disruptions, such as traffic accidents, social events, and adversarial weather conditions, often occur unexpectedly in real-world networks, examining the robustness, resilience, reliability, and even safety of transportation systems is highly crucial (Ganin et al., 2019; Zayandehroodi et al., 2025). What makes it even more challenging is the continuous growth of motorized traffic. For instance, an approximate 50% increase in traffic demand can be observed in both the U.S. (U.S. Department of Transportation, 2019) and Switzerland (Federal Statistical Office, 2020) over the past few decades. Such growth can lead to not only an escalation of congestion but also more traffic accidents (Albalate and Fageda, 2021). To picture the consequences, recent work in Sun et al. (2024) has demonstrated the fragile nature of road transportation networks with mathematical proof, demonstrating that performance degrades exponentially with linearly increasing magnitudes of disruptions. Therefore, transportation systems must be designed to secure a decent level of service even when faced with unexpected disruptions of unforeseen magnitudes and the exponentially growing negative effects.
@@
To address such issues, the concept of antifragility has shed light on a potential solution. Echoing the sentiment of the well-known maxim What does not kill me makes me stronger in Nietzsche (1889), antifragility was first introduced in the reputable general-audience publication Antifragile: Things That Gain from Disorder (Taleb, 2012). It has then been mathematically formulated as an academic concept in Taleb and Douady (2013), providing insights into designing systems that can benefit from disruptions and perform better under growing volatility and randomness. Its counterpart concept, fragility, can be dated in complexity theory in Vespignani (2010), indicating a cascading effect of interdependent variables in complex networks, such as in transportation networks (Cats and Hijner, 2021). Ever since the concept was proposed, antifragility has gained substantial interest from both the public and academia, particularly in the complex systems and risk engineering communities (Aven and Krohn, 2014; Aven, 2015; Thekdi and Aven, 2019; Aven, 2022; Axenie et al., 2024). However, the design principles for realizing antifragile transportation systems and control are generally unexplored. One promising approach to induce antifragility is to leverage learning-based algorithms, such as Reinforcement Learning (RL), since RL agents can gradually adjust their decision-making when deployed to an environment subject to variations.
The primary goal of this paper is to design an antifragile perimeter control algorithm capable of adapting to increasing magnitudes of disruptions, reflecting the effects of urban densification and rising traffic demand. Fig. 1 presents a map of contributions while highlighting the key references. The contributions of this research can be summarized as follows:
  1. 1.
    Conceptual contribution: We distinguish the concept of antifragility as opposed to other related and commonly used terms in the risk engineering and transportation domains.
  2. 2.
    Methodological contribution: We introduce how antifragility can be incorporated into RL algorithms to achieve superior performance compared to benchmark methods, tackling both fragile performance and observability issues.
  3. 3.
    Evaluatory contribution: We adopt a skewness-based quantitative indicator to showcase the antifragile properties of our proposed algorithm under increasing disruptions.
  4. 4.
    General applicability: We further tested the effectiveness of the proposed antifragile module on other state-of-the-art RL algorithms and validated its applicability.
@@ 2 Literature review
This section reviews the relevant literature on three topics intertwined in this work. First, a macroscopic traffic model and the control strategy applied in this work are introduced, which serve as the basis of model dynamics for the simulation environment. Next, we review how robustness or resilience can be induced with RL algorithms to counter real-world model uncertainty. Finally, related terminologies to antifragility are distinguished while introducing the design philosophy of an antifragile system.
@@ 2.1 Macroscopic fundamental diagram and perimeter control
Alleviating urban network congestion can be realized through various traffic control strategies. Since Geroliminis and Daganzo (2008) demonstrated the presence of the Macroscopic Fundamental Diagram (MFD) with empirical data, the mathematical relationship between flow-density-speed has paved the path for the development of control strategies at a macroscopic level, enabling more computationally feasible real-time traffic control strategies for large-scale networks, such as perimeter control (Keyvan-Ekbatani et al., 2012; Geroliminis et al., 2013; Kouvelas et al., 2017), congestion pricing (Genser and Kouvelas, 2022), route guidance (Fu et al., 2022).
@@
Perimeter control is among the strategies that have attracted immense attention and research efforts. By mitigating incoming flows from adjacent regions into a protected zone, the traffic density in the protected area remains below the critical density to uphold a satisfactory serviceability. Geroliminis et al. (2013) proposed an optimal perimeter control method using Model Predictive Control (MPC) and proved its effectiveness compared to a greedy controller in a cordon network. Later, Sirmatel and Geroliminis (2020) introduced a Moving Horizon Estimation (MHE) scheme together with MPC to tackle further measurement noise and the observability issues. One major issue of the previous works is the assumption of MFD homogeneity, and a substantial amount of effort has been made in investigating the partitioning algorithms so that a well-defined MFD can be created for each sub-network, as in Ambühl et al. (2019), creating a multi-reservoir system for applying perimeter control (Aboudolas and Geroliminis, 2013). However, MFDs in the real world can hardly be well-defined, as demonstrated in Ambühl et al. (2021) with loop detector data over a year. Wang et al. (2015) also showed that adverse weather conditions and traffic incidents can alter the shape of the MFDs, and even a recovery from the peak-hour congestion may lead to a hysteresis (Gayah and Daganzo, 2011). These phenomena could potentially violate the mathematical model that serves as the foundation for the established model-based perimeter controllers. Therefore, Zhou and Gayah (2021) introduced an RL-based algorithm that exhibits superior performance and offers the advantage of being model-free.
2.2 Leveraging RL to induce robustness, resilience, and reliability
Similar to traffic networks and system control, real-world systems of all kinds can be subject to disruptions. As RL offers flexibility in defining state representations, actions, and rewards, researchers have leveraged such features to enhance system performance under adversarial events. Here, we present a review of studies that apply RL in reliability engineering and traffic engineering domains, claiming enhancements in robustness, resilience, reliability, or related properties.
Most existing studies on data-driven maintenance scheduling algorithms for infrastructure systems or mechanical equipment prioritize costs or efficiency, often lacking an orientation to reliability or resilience. To bridge such gaps, Yuan et al. (2025) proposed an innovative reliability reward component based on the Gamma-Wiener process, enabling the RL agent to achieve a balance between dynamic reliability and maintenance cost of mechanical equipment. Likewise, Liang et al. (2026) proposed a specialized resilience metric, which encourages the RL agent to accelerate post-disaster recovery while explicitly considering system resilience, while Yin et al. (2025) couples the DistFlow model for feasible resilience assessment and Bayesian learning for the quantification of system uncertainties to promote resilient post-disruption recovery. Risk and uncertainty can also be directly modeled in the reward function as in Chiou (2024). On top of reward shaping, Xu et al. (2026) also introduced a phase-adaptive discount factor to enhance the resilience of combat system-of-systems.
In traffic networks, RL algorithms can also be engineered to strengthen the robustness or resilience of different control systems, such as traffic light control or perimeter control. For instance, Rodrigues and Azevedo (2019) induced robustness by adding the elapsed time since the last green signal for each phase, Tan et al. (2020) experimented with speed or residual queue as an additional state representation in the RL algorithm, Chu et al. (2020) supplemented the control policies of neighboring intersections as additional information to the agents, and Zhou and Gayah (2023) used an extra binary congestion indicator in the state space. As a result, the algorithms are provided with additional information related to disturbances or disruptions.
@@ 2.2 Leveraging RL to induce robustness, resilience, and reliability
The analysis of the state-of-the-art RL studies considers reversals or sudden changes in the state-action-reward dynamics, which evoke unanticipated uncertainty. The problem in these contexts is often to respond to unexpected results appropriately, since they might indicate a shift in the environment. In this case, exploration refers to the process of searching for new information to improve the RL agent’s understanding of the traffic dynamics under disruptions, which would then be used to identify better courses of action. By exploring how robustness, resilience, and reliability can be achieved using RL-based traffic control algorithms, we can potentially develop antifragile traffic control systems through similar approaches.
@@ 2.3 Antifragility: definitions and distinctions
Before diving into antifragile system design, we first need to review the distinctions between antifragility and other closely related terms, which include robustness, resilience, reliability, stability, and adaptiveness. We follow the definitions proposed in Zhou et al. (2019), that robustness is concerned with assessing a system’s capacity to preserve its initial state and resist performance deterioration under minor disturbances, while resilience emphasizes the ability and speed of a system to recover from major disruptions to the original state. Reliability in transportation has a much broader range of meanings. Pennetti et al. (2020) proposed travel time reliability, which has a similar flavor to robustness in measuring deviation from normal operation, but focuses on the probability of such deviation exceeding a certain threshold. Cats et al. (2017) introduced a reliability assessment indicator to account for performance loss for the entire range of possible capacity reductions, whereas robustness pertains merely to a single disruption event. Stability can also bear multiple definitions based on the context, such as describing the closed-loop stability of system control as in Sirmatel and Geroliminis (2021) or MFd hysteresis on the traffic flow theory basis (Gayah and Daganzo, 2011). Adaptiveness describes the ability of systems to alter their traits to satisfy autonomy in different environments (Hooker, 2011), which aligns with the concept of proto-antifragility proposed in Taleb (2012). However, adaptiveness does not necessarily guarantee a consistent performance improvement under growing magnitudes of disruptions. Therefore, Taleb (2012) introduced the concept of antifragility, emphasizing the concave response of the system under increasing disruptions, which can be mathematically formulated with Jensen’s inequality $\mathbb{E}[g(X)]\leq g(\mathbb{E}[X])$. A graphical comparison between robustness, resilience, adaptiveness, and antifragility is illustrated and described in Fig. 12. It should be noted that while (anti-)fragility can be an innate property of a system, proper intervention and control strategies can lead the system to be more antifragile (Axenie et al., 2024). For instance, building on the mathematical proof by Sun et al. (2024) that urban road networks are naturally fragile, this paper seeks to develop antifragile solutions that mitigate such fragility through perimeter control.
@@
Ever since the concept of antifragility was proposed, it has become an increasingly popular concept across various disciplines in recent years, such as energy (Coppitters and Contino, 2023), electricity (Rachunok and Nateghi, 2020), biology (Kim et al., 2020), medicine (Axenie et al., 2022), cyber-systems (Chatterjee and Thekdi, 2020), robotics (Axenie and Saveriano, 2023), and machine learning (Pravin et al., 2024). Researchers have also proposed methods to incentivize the antifragile property of a system by emphasizing the derivatives of system states to capture the temporal evolution patterns of the system dynamics, i.e., how fast the system state deviates toward possible black swan events and the curvature of this deviation (Taleb and Douady, 2013; Axenie et al., 2022). With this additional information, the system can potentially anticipate ongoing disruptions and become more proactive in facing drastic changes. Similar to its function in resilience (Kamalahmadi et al., 2022), redundancy can also be added to the system to induce antifragility Munoz et al. (2022). Other feasible approaches also include time-scale separation and attractor dynamics (Axenie et al., 2022). However, leveraging the potential of antifragile system design for the operation of transportation networks is still an unexplored notion.
@@ 3 Problem formulation
This paper studies perimeter control between two homogeneous cordon-shaped urban networks, as in Geroliminis et al. (2013), with the inner region representing a city center, as shown in Fig. 2(a). The total number of vehicles in region $i$ at time $t$ is denoted as $n_{i}(t)$, while the Origin-Destination (OD) from region $i$ to region $j$ can be further divided as $n_{ij}(t)$. The inner and outer regions are assumed to have different MFDs to represent the capacity difference of accommodating vehicles within the network between the city center and the surrounding region, which are defined as $G_{i}(n_{i}(t))$ as illustrated in Fig. 2(b). The total trip completion rate $M_{i}(t)$ for region $i$ at time $t$ can be determined through the corresponding MFD, which comprises both the intraregional trip completion $M_{ii}(t)$ and the interregional transfer flow $M_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. While $q_{i}(t)$ denotes the new demand within region $i$ at time $t$, $q_{ij}(t)$ represents the additional demand based on the OD. To protect a certain region from being overflown by high traffic demand, the percentage of the transfer flow $M_{ij}(t)$ allowed to get across the perimeter is regulated by the perimeter controllers, denoted as $u_{ij}(t)\>(i\neq j)$ with $i,j\in\{1,2\}$. All notations applied in this work, including those in the following sections, are summarized in Table 1.
@@
1. Notations for the problem formulation (continuous form)
$t$ Current time